External Funds Needed 3.13 Using the definitions below show that EFN can be written as pHist Asset needs will equal A x & The addition to retained earnings will equal PM(S)b x (1+g S Previous years sales A Total assets D Total debt E Total equity g Projected growth in sales PM Profit margin b- Retention (plowback) ratio MINICASE Ratios and Financial Planning at East Coast Yachts Dan Ervin was recently hired by East Coast Yachts individuals for pleasure use. Occasionally, a yacht to assist the company with its short-term financial is manufactured for purchase by a company for planning and also to evaluate the companys finan- business purposes. cial performance. Dan graduated from university five years ago with a finance degree, and he has a number of manufacturers. As with any industry been employed in the treasury department of a For- there are market le tune 500 company since then. The custom yacht industry is fragmented. with the industry ensures that no manufacturer domi- East Coast Yachts was founded 10 years ago nates the market. The competition in the market as by Larissa Warren. The companys operations well as the product cost, ensures that attention to are located near Lunenburg Nova Scotia, and the detail is a necessity. For instance. East Coast Yachts company is structured as an LLC (limited liability will spend 80 to 100 hours on hand buffing the company. The company has manufactured mid- stainless steel stem-iron, which is the metal cap on custom high-performance yachts for clients the yachts bow that conceivably could collide with a over this period. and its products have received dock or another boat. excellent reviews for safety and reliability. The To get Dan started with his analysis, Larissa has companys yachts have also recently received provided the following financial statements. Dan the highest award for customer satisfaction. has gathered the industry ratios for the yacht manu- The yachts are primarily purchased by wealthy facturing industry East Coast Yachts 2015 Statement of Comprehensive Income $234,300,000 165,074,000 27,991,000 7,644,000 33,591,000 Cost of goods sold Earnings before interest and taxes s 29,378,400 11,751,360 S 17.627.040 Taxable income 5,288,112 Addition to retained earnings
media%2Fc9f%2Fc9fb1ee3-569f-428c-bdd5-1e
Answer #1

1. Calculation of all the ratios listed in the Industry table for East Coast Yachts:-

Amount in $

Current Ratio- Current Assets/ Current Liabilities

17,582,000/23,689,300 = 0.7422

Quick Ratio : Current assets - Inventory/Current Liabilities

17,582,300 - 7,363,700/23,689,300 = 0.4314

Total asset turnover : Sales / Total assets

234,300,000 / 130,338,900 = 1.7976

Inventory Turnover = Cost of goods sold / Inventory

165,074,000 / 7,363,700 = 22.4172

Receivables Turnover = Sales / Accounts receivable

234,300,000 / 6,567,600 = 35.6751

Debt Ratio = Total assets - Total equity / Total assets

130,338,900 - 66,169,600 / 130,338,900 = 0.4923

Debt to equity ratio = Total debt / Total equity

64,169,300 / 66,169,600 = 0.9698

Equity Multiplier = Total assets / Total equity

130,338,900 / 66,169,600 = 1.9698

Interest Coverage = EBIT / Interest

33,591,000 / 4,212,600 = 7.9740

Profit margin = Net Income / Sales

17,627,040 / 234,300,000 = 7.52%

Return on assets = Net income / Total assets

17,627,040 / 130,338,900 = 13.52%

Return on equity = Net income / Total equity

17,627,040 / 66,169,600 = 26.64%

-----------------------------------------------------------------------------------------------------------------------------------------------------

2. Performance comparison with Industry:  

Liquidity ratios:

East Coast Yachts (ECY) has 0.74 in current assets for every $1 in current liabilities. Current ratio less than 1 means that the net working capital of firm is negative. This shows that company has debt taking reserve. ECY has low current ratio as it is below the Industry median.

Quick ratio is 0.4314 which shows that they use lot of cash to purchase inventory. This is above the Industry median which can allow ECY to purchase more inventory.

Total asset turnover ratio is 1.7976. This is above the Industry quartile stating that ECY makes profitable sales. This shows that for every $1 in assets ECY generated $1.7976 in sales.

Other turnover ratios (Inventory turnover and Receivables turnover) are above the Industry quartile. This means that ECY is more efficient in the industry in using its assets to generate sales. So this is positive for ECY.

Financial leverage ratios (Debt ratio, debt to equity ratio, equity multiplier, interest coverage) all are at median but above the lower quartile. It has less debt as compared to other companies in the industry and is in the normal range.

Profit margin of the company is above the Industry median, Return on assets is little bit higher than the Indusrty median and Return on equity is far higher than the Industry median. The company seems to perform well and good in profitability.

Finally, the overall performance of East Coast Yachts seems good.

------------------------------------------------------------------------------------------------------------------------------------

3. Sustainable growth rate: -

ROE * (1-Dividend Payout Ratio)

Dividend Payout = Total Dividend / NEt Income

5288112/17627040

= 0.3

ROE is 26.64% or 0.2664

Sustainable growth rate is = 26.64% * (1-0.3)

= 18.65%